For the following overall cell reaction, write the oxidation and reduction half-reactions, calculate the standard cell potential, E^0cell and determine if the reaction is spontaneous or not. Show your work.


Fe^3+(aq) + Cu^+(aq) —> Fe^2+(aq) + Cu^2+(aq)


1
Expert's answer
2022-06-05T22:51:03-0400

1.

Fe3+ + e- => Fe2+ — reduction half-reaction

Cu+ - e- => Cu2+ — oxidation half-reaction

2.

Fe3+ + 1e- => Fe2+ E0 = 0,771 V

Cu+ - 1e- => Cu2+ E0 = -0,16 V

E0cell = 0,771 + (-0,16) = 0,611 V

If E°cell > 0, then the process is spontaneous (galvanic cell) If E°cell < 0, then the process is non-spontaneous (the voltage must be supplied, as in an electrolytic cell).

therefore, the reaction goes by spontaneous